### v1

# neurips_2025.bib Edge-ASR: Towards Low-Bit Quantization of Automatic Speech Recognition Models

Abstract
Recent advances in Automatic Speech Recognition (ASR) have demonstrated remarkable accuracy and robustness in diverse audio applications, such as live transcription and voice command processing. However, deploying these models on resource-constrained edge devices (e.g., IoT device, wearables) still presents substantial challenges due to strict limits on memory, compute and power. Quantization, particularly Post-Training Quantization (PTQ), offers an effective way to reduce model size and inference cost without retraining. Despite its importance, the performance implications of various advanced quantization methods and bit-width configurations on ASR models remain unclear. In this work, we present a comprehensive benchmark of eight state-of-the-art (SOTA) PTQ methods applied to two leading edge-ASR model families, Whisper and Moonshine. We systematically evaluate model performances (i.e., accuracy, memory I/O and bit operations) across seven diverse datasets from the open ASR leaderboard, analyzing the impact of quantization and various configurations on both weights and activations. Built on an extension of the LLM compression toolkit, our framework integrates edge-ASR models, diverse advanced quantization algorithms, a unified calibration and evaluation data pipeline, and detailed analysis tools. Our results characterize the trade-offs between efficiency and accuracy, demonstrating that even -bit quantization can succeed on high capacity models when using advanced PTQ techniques. These findings provide valuable insights for optimizing ASR models on low-power, always-on edge devices.
1 Introduction
Real-time Automatic Speech Recognition (ASR) models have revolutionized the natural audio processing pipeline and achieved remarkable accuracy and robustness in diverse applications such as live transcription, voice command recognition and assistive use cases ([ASRSurvey]). As interest in on-device ASR continues to grow, there is an emerging demand for models that operate efficiently on low-power, always-on edge hardware (e.g., IoT device, wearables). Among the state-of-the-art (SOTA) edge-ASR models, OpenAI’s Whisper ([Whisper]) has demonstrated strong accuracy in compact model sizes, making it a practical candidate for edge inference. Meanwhile, Useful Sensors’ Moonshine ([Moonshine]) reduces latency and adapts well to variable-length audio inputs, further optimized for resource-constrained deployment. Both model families currently lead the open ASR leaderboard ([Leaderboard]) across seven evaluation datasets, among model sizes ranging from M to M parameters.
Despite the advances, deploying ASR models on edge devices remains challenging due to memory, compute and power constraints. Most existing low-power neural processing units (NPUs), microcontrollers and digital signal processors (DSPs) are designed as fixed-point inference accelerators, making full precision models impractical. Quantization, particularly Post-Training Quantization (PTQ), addresses this by mapping weights and activations to low-bit data formats without retraining, directly reducing model size, latency and power consumption while maintaining acceptable accuracy. Its simplicity and hardware compatibility have driven widespread industry adoption.
While many advanced PTQ techniques have been proposed and studied in the context of large language models (LLMs) ([QuantLLM]), their applications to edge-ASR remain unexplored. Exisiting quantization toolkits and benchmarks ([LLMC]) focus on LLM model architectures, and task-specific datasets that do not apply to speech recognition directly. Furthermore, although the open ASR leaderboard provides standardized full precision baselines, it lacks coverage of quantized models, leaving the performance trade-offs from quantization in this domain largely uncharacterized.
In this work, we fill this gap by benchmarking eight PTQ methods applied to two Transformer-based edge-ASR families, Whisper and Moonshine. The quantization methods are categorized into three primary strategies: transformation-based, reconstruction-based and rounding optimization. We systematically evaluate quantization of weights and activations across seven public ASR datasets, measuring accuracy, memory I/O and bit operations (BOPs) ([BOP]). We also perform ablation studies on calibration data sources and sizes, and analyze quantization sensitivity. All experiments are built on an extension of the LLM compression toolkit ([LLMC]), enabling a reproducible and extensible ASR quantization framework. The overall workflow is demonstrated in Figure 1.
Contributions. (a) We deliver the first systematic benchmark of quantized edge-ASR models, assessing eight advanced PTQ methods over multiple datasets, architectures and configurations, to establish a quantized standard for the open ASR leaderboard. (b) We extend and contribute to the LLM compression toolkit to support ASR models, enabling an end-to-end, extensible quantization workflow for speech recognition. (c) Through extensive evaluations, analysis and comprehensive comparisons from various factors (e.g., model accuracy, memory I/O, BOPs), we provide valuable insights into how different quantization methods and configurations impact model performance, offering guidance for deploying ASR on memory and compute constrained low-power edge devices.
2 Preliminaries
2.1 Quantization Algorithms
Quantization maps high-precision floating-point values to discrete low-precision levels, reducing memory footprint and accelerating inference. We focus on integer-uniform Post-Training Quantization, which is broadly supported by edge hardware accelerators. The general quantization formula for weights or activations is given by:
where denotes the floating-point value, denotes the low-precision quantized value. and denote the scaling factor and zero-point, respectively. represents for the quantized range based on the target bit-width . For symmetric quantization, the zero-point is set to zero.
Post-Training Quantization algorithms can be broadly categorized into three primary strategies: scaling-based transformation, reconstruction-based optimization, and rounding-based optimization. Scaling-based transformation techniques, such as SmoothQuant ([SmoothQuant]), AWQ ([AWQ]), and OmniQuant ([OmniQuant]), apply learned or heuristic scaling to weights and activations, reshaping their distributions to be more quantization-friendly. Reconstruction-based methods, such as GPTQ ([GPTQ]), reconstruct block outputs by minimizing quantization error iteratively, preserving the original model behavior after quantization. Rounding-based optimizations refine the rounding process to better preserve accuracy. Representitive methods are naïve round-to-nearest (RTN) and TesseraQ ([TesseraQ]), where an iterative and progressive adaptive rounding is proposed. In addition, methods like QUIK ([QUIK]) and SpQR ([SpQR]) use hybrid quantization strategies to balanace model accuracy and overall compression ratio. Table 1 summarizes the eight evaluated quantization algorithms.
| Category | Algorithm | Strategy |
| Scaling-based Transformation | SmoothQuant | |
| AWQ | ||
| OmniQuant | , and | |
| Reconstruction-based | GPTQ | |
| Rounding Optimization | RTN | |
| TesseraQ | ||
| Hybrid | QUIK, SpQR |
2.2 edge-ASR Models
To evaluate the impact of post-training quantization techniques on edge-ASR performance, we select two state-of-the-art open-source ASR model families: OpenAI’s Whisper ([Whisper]) and Useful Sensors’ Moonshine ([Moonshine]). Both model families lead the open ASR leaderboard across seven diverse datasets and feature compact model sizes (M to M parameters), making them well-suited for deployment on low-power, always-on edge devices.
Whisper adopts an encoder-decoder Transformer architecture. Input audio is resampled to kHz, segmented into -second chunks, and converted into channel log-Mel spectrograms for the encoder. The decoder generates text, intermixed with special tokens for tasks such as language identification, multilingual speech transcription, and English translation. Whisper’s large-scale pretraining yields strong robustness in real-world scenarios.
Moonshine follows a similar encoder-decoder Transformer structure, but is explicitly optimized for low-latency on-device inference. It takes raw audio input and incorporates rotary positional embedding (RoPE), allowing support for variable-length audio inputs. Compared to Whisper, Moonshine achieves lower latency on short segments while maintaining competitive accuracy, making it an attractive choice for constrained edge environments.
Together, Whisper and Moonshine represent significant advances in edge-ASR models, addressing accuracy, model capacity and latency challenges. Detailed architecture specifications for each model family are provided in Appendix A.
2.3 Benchmarks and Datasets
Following the open ASR leaderboard, we benchmark eight PTQ methods on two edge-ASR model families across seven evaluation datasets: AMI corpus ([AMI]), Eearnings-22 ([Earnings-22]), GigaSpeech ([GigaSpeech]), LibriSpeech ([Librispeech]), SPGISpeech ([SPGISpeech]), TED-Lium ([TED-LIUM]) and VoxPopuli ([VoxPopuli]). These datasets cover a diverse range of domains, accents and acoustic conditions, providing a robust performance baseline.
For each quantization method, we explore multiple configurations: varying weight and activation bit-widths and quantization granularities. To align with typical support in low-power inference accelerators, we evaluate per-channel and per-group quantization for weights, and per-tensor and per-token quantization for activations. Our evaluation reports a full spectrum of deployment-relevant metrics, including word-error-rate (WER), model size, memory I/O and BOPs. The comparisons offer end-to-end insights into the trade-offs between accuracy and efficiency under different quantization strategies.
2.4 Statistical Analysis
To better understand the source of quantization degradation in specific layers or configurations, we incorporate a statistical analysis based on Kurtosis value of a tensor, defined as:
where is the number of data points, is the mean and is the standard deviation. Intuitively, the kurtosis value quantifies the outlier condition of a certain distribution. A higher kurtosis indicates a distribution with heavy tails and outliers, which may result in larger quantization errors. Conversely, a low kurtosis suggests a distribution with fewer outliers and typically lower quantization loss. By analyzing layer-wise kurtosis for both weights and activations, we can identify which components are most vulnerable to quantization and may benefit from techniques such as mixed-precision.
3 Benchmarking Quantized edge-ASR Models
This section presents a comprehensive benchmarking of PTQ methods on edge-ASR models. We begin with the experimental setup (subsection 3.1), followed by an in-depth quantization result and analysis (subsection 3.2), covering bit-width selection, ultra-low bit quantization, granularity, weight clipping, and the use of symmetric versus asymmetric quantization. We then compare deployment-relevant metrics, including WER, model size, memory I/O and BOPs in subsection 3.3. Additional ablation studies on the impact of calibration data source and size are further discussed in Appendix E.
3.1 Experimental Setup
We outline the experimental setup used to evaluate quantized edge-ASR models. Additional implementation details are available in Appendix A - C.
Models. We assess two ASR families optimized for edge deployment. Three variants: Tiny (M), Base (M) and Small (M), are selected from the Whisper family, while the Moonshine family includes Tiny (M) and Base (M). Both model families share an encoder-decoder Transformer architecture but differ in positional embedding and configurations. We employ a byte-level BPE tokenizer (vocabularies of for Whisper and for Moonshine), and apply standard text normalization before computing WER.
Datasets. Following the open ASR leaderboad, we evaluate on seven benchmark datasets. These datasets cover diverse domains, accents and speaking styles, with audio durations ranging from to hours. Our floating-point baselines match reported leaderboard results ([Leaderboard]). For algorithms that require calibration, we use a fixed set of English utterances sampled from the Mozilla Common Voice ([CommonVoice]) that are different from any evaluation dataset, ensuring a fair comparison. Table 2 summarizes the full benchmark setup.
| PTQ algorithms | SmoothQuant, AWQ, OmniQuant, GPTQ, RTN, TesseraQ, QUIK, SpQR |
| Granularities | weights: per-channel, per-group |
| activations: per-tensor, per-token | |
| edge-ASR models | Whisper (Tiny, Base, Small), Moonshine (Tiny, Base) |
| Datasets | AMI corpus, Eearnings-22, GigaSpeech, LibriSpeech (clean, other), SPGISpeech, TED-Lium, VoxPopuli |
| Evaluation metrics | WER, kurtosis value, model size, memory I/O, bit operations (BOPs) |
3.2 Quantization Analysis
3.2.1 Bit-width selection
To determine the optimal bit-width for ASR models, we evaluate WER under various PTQ algroithms across all benchmark datasets. Table 3 reports the averaged WER of the quantized edge-ASR models across datasets, alongside its floating-point baseline. Detailed per-dataset WER performances of each model are provided in Appendix D.
Our results show that PTQ with -bit weights and -bit activations (w8-a8) reliably preserves model performance across all models and quantization algorithms. Within both model families, reducing weight precision to bits leads to notable degradation in smaller models such as Whisper Tiny (Table 11) and Moonshine Tiny (Table 14), while larger models, including Whisper Base (Table 12), Whisper Small (Table 13) and Moonshine Base (Table 15) exhibit stronger robustness under the same low-bit quantization. For the Moonshine family, SmoothQuant performs poorly at low weight bit-width, and only becomes effective when weight precision increases to -bit.
| Bits | Method | averaged WER % | ||||
| Whisper Tiny | Whisper Base | Whisper Small | Moonshine Tiny | Moonshine Base | ||
| Float | Baseline | 12.80 | 10.32 | 8.59 | 12.72 | 9.99 |
| w4-a8 | SmoothQuant | 31.78 | 16.81 | 9.68 | 72.99 | 77.22 |
| AWQ | 21.52 | 11.87 | 8.82 | 15.32 | 10.98 | |
| OmniQuant | 20.31 | 11.77 | 8.64 | 14.74 | 10.74 | |
| GPTQ | 18.57 | 11.43 | 8.74 | 14.15 | 10.38 | |
| RTN | 21.30 | 11.44 | 8.78 | 25.38 | 13.48 | |
| TesseraQ | 17.05 | 11.40 | 8.73 | 17.95 | 12.08 | |
| QUIK | 14.69 | 11.06 | 8.86 | 14.32 | 10.37 | |
| SpQR | 15.75 | 11.14 | 8.60 | 14.15 | 10.31 | |
| w4-a16 | SmoothQuant | 21.90 | 12.49 | 9.00 | 69.81 | 14.20 |
| AWQ | 17.83 | 11.41 | 8.92 | 15.12 | 10.57 | |
| OmniQuant | 16.85 | 11.59 | 8.92 | 14.47 | 10.54 | |
| GPTQ | 16.46 | 11.16 | 8.77 | 13.61 | 10.28 | |
| RTN | 17.78 | 11.36 | 8.86 | 23.23 | 12.94 | |
| TesseraQ | 17.58 | 11.42 | 8.87 | 15.21 | 10.78 | |
| QUIK | 14.73 | 11.30 | 8.90 | 14.24 | 10.40 | |
| SpQR | 14.24 | 10.94 | 8.65 | 13.81 | 10.14 | |
| w8-a8 | SmoothQuant | 13.93 | 10.82 | 8.55 | 13.42 | 38.87 |
| AWQ | 14.04 | 10.40 | 8.64 | 13.12 | 10.16 | |
| OmniQuant | 13.86 | 11.26 | 8.44 | 13.04 | 10.14 | |
| GPTQ | 13.65 | 10.52 | 8.45 | 13.09 | 10.15 | |
| RTN | 13.39 | 10.51 | 8.56 | 13.05 | 10.09 | |
| TesseraQ | 13.16 | 10.59 | 8.47 | 13.50 | 10.29 | |
| QUIK | 12.94 | 10.18 | 8.63 | 12.81 | 10.04 | |
| SpQR | 14.04 | 10.54 | 8.54 | 13.04 | 10.12 | |
| w8-a16 | SmoothQuant | 12.76 | 10.49 | 8.62 | 12.78 | 9.98 |
| AWQ | 12.85 | 10.21 | 8.49 | 12.78 | 10.06 | |
| OmniQuant | 13.09 | 10.41 | 8.58 | 12.73 | 10.05 | |
| GPTQ | 12.91 | 10.40 | 8.61 | 12.77 | 9.99 | |
| RTN | 12.78 | 10.30 | 8.64 | 12.76 | 10.03 | |
| TesseraQ | 12.68 | 10.17 | 8.64 | 12.74 | 10.05 | |
| QUIK | 12.72 | 10.28 | 8.62 | 12.74 | 10.04 | |
| SpQR | 12.73 | 10.25 | 8.59 | 12.76 | 10.03 | |
Interestingly, the w8-a8 configuration consistently surpasses w4-a16, highlighting the importance of weight quantization. Overall, reconstruction-based and hybrid quantization methods outperform scale-based transformations under low-bit settings. These patterns are consistent across all Whisper and Moonshine variants, confirming that while -bit precision is generally safe for deployment, lower-bit weight quantization requires careful algorithmic design and model capacity considerations. As a reference, the runtime cost of each quantization algorithm is available in Appendix D.6.
3.2.2 Ultra-low bit quantization
To explore the limits of PTQ under extreme compression, we evaluate ultra-low bit weight quantization on the Whisper Base and Moonshine Base models. In these experiments, weights are quantized asymmetrically, and activations use per-token quantization to maximize their representation power.
Model Method w3-a16 (WER %) AMI Earnings-22 GigaSpeech Libri clean Libri other SPGISpeech TED-Lium Voxpopuli Avg. Whisper Base Float 21.13 15.09 12.87 4.28 10.36 4.27 4.85 9.75 10.32 SmoothQuant 100.94 207.28 113.70 199.29 185.28 128.14 144.16 216.92 161.96 AWQ 31.10 36.80 19.71 9.96 21.65 11.15 9.58 22.29 20.28 OmniQuant 30.79 26.64 17.85 7.91 15.24 8.55 7.70 15.74 16.30 GPTQ 36.65 35.24 24.93 11.38 22.27 10.98 11.48 22.15 21.89 RTN 27.23 29.85 17.49 9.19 15.27 8.78 7.38 16.28 16.43 TesseraQ 28.20 27.36 16.88 8.32 16.81 8.64 7.13 16.05 16.17 QUIK 23.16 20.05 14.28 6.36 13.67 5.95 5.63 14.18 12.91 SpQR 22.64 18.36 14.44 5.01 12.73 5.18 6.50 12.34 12.15 Moonshine Base Float 17.07 17.69 12.11 3.26 8.28 5.46 5.24 10.79 9.99 SmoothQuant 33.95 36.94 21.78 13.74 29.93 16.69 12.41 22.42 23.48 AWQ 20.86 24.37 14.12 4.24 11.25 7.05 6.49 12.36 12.59 OmniQuant 21.21 22.85 13.91 4.22 10.72 7.34 6.41 12.90 12.45 GPTQ 18.92 20.18 13.08 3.92 9.79 6.48 5.96 12.10 11.30 RTN 90.01 117.00 74.32 70.87 101.09 89.60 73.99 115.73 91.58 TesseraQ 93.03 122.89 76.53 70.73 105.21 94.94 77.54 119.75 95.08 QUIK 20.44 20.12 13.21 3.95 10.23 6.95 6.01 11.72 11.58 SpQR 18.45 19.33 12.49 3.75 9.11 5.89 5.13 11.23 10.67
Enforcing -bit weight precision causes all PTQ methods to fail, resulting unacceptably high WER. For weights below -bit, Quantization-Awared Training (QAT) becomes essential. However, at -bit weight precision, several algorithms produce promising results. As shown in Table 4, OmniQuant and hybrid strategies (e.g., QUIK, SpQR that preserves higher precision for the top outliers) maintain reasonable accuracy. In contrast, scaling-based transformation such as SmoothQuant perform poorly on the Whisper Base model. Similarly, RTN and TesseraQ struggle on the Moonshine Base model at this precision, indicating that their rounding-based strategies alone are insufficient at ultra-low precision.
To further examine each algorithm’s internal behavior, we analyze the Whisper Base model and present the averaged kurtosis value of weights and activations over all layers in Figure 2. This provides a statistical perspective on the distribution characteristics after quantization. We observed that scaling-based transformation methods (SmoothQuant, AWQ) achieve lower kurtosis for activations but at the cost of increased weight kurtosis, due to the design principle of migrating the quantization difficulty from activations to weights. This explains significant performance degradation under ultra-low bit settings, even when activations remain at bits. These results underscore the need for algorithmic roubustness and outlier handling when pushing PTQ to its limits.
3.2.3 Quantization granularity
To assess the impact of weight and activation quantization granulairty, we select three representitive PTQ algorithms (AWQ, GPTQ and TesseraQ), one from each categrory, and apply -bit uniform quantization to weights under two schemes: per-channel and per-group. -bit per-tensor and per-token quantization are applied on activations. Table 5 reports the averaged WER across all seven datasets for each model and granularity setting.
| Method | granularity | w4-a16 Models (avg. WER %) | |||||
| weights | activations | Whisper tiny | Whisper base | Whisper small | Moonshine tiny | Moonshine base | |
| Float | 12.80 | 10.32 | 8.59 | 12.72 | 9.99 | ||
| AWQ | per-channel | per-tensor | 64.34 | 85.64 | 9.94 | 15.67 | 11.16 |
| GPTQ | 28.04 | 15.32 | 9.27 | 14.60 | 10.74 | ||
| TesseraQ | 43.58 | 18.06 | 9.42 | 16.26 | 11.43 | ||
| AWQ | per-group, coarse | per-tensor | 23.26 | 12.64 | 8.86 | 15.17 | 10.79 |
| GPTQ | 16.84 | 11.60 | 8.81 | 13.91 | 10.50 | ||
| TesseraQ | 20.51 | 11.75 | 8.96 | 15.65 | 11.20 | ||
| AWQ | per-group, fine | per-tensor | 17.83 | 11.41 | 8.92 | 15.12 | 10.57 |
| GPTQ | 16.46 | 11.16 | 8.77 | 13.61 | 10.28 | ||
| TesseraQ | 17.58 | 11.42 | 8.87 | 15.21 | 10.78 | ||
| AWQ | per-group, fine | per-token | 17.82 | 11.38 | 8.96 | 14.95 | 10.64 |
| GPTQ | 16.72 | 11.88 | 8.95 | 13.67 | 10.30 | ||
| TesseraQ | 17.78 | 11.36 | 8.86 | 15.24 | 10.77 | ||
Our results show that GPTQ consistently delivers the lowest averaged WER across all weight granularity configurations, demonstrating the strong robustness to coarse quantization. For larger models (Moonshine Base and Whisper Small), per-channel weight quantization shows reasonable accuracy, making it an efficient option. In contrast, smaller models (Whisper Tiny and Whisper Base) benefit notably from per-group quantization, where finer-grained scaling helps to mitigate quantization noise and significantly improves WER. We also observe that activation granulairty has minimal impact at -bit precision. These findings highlight the importance of choosing appropriate weight quantization granularity, especially when targeting low capacity models on hardwares with limited quantization support.
3.2.4 Weight clipping
The technique of weight clipping is commonly used to constrain the range of weight values prior to quantization. By suppressing extreme outliers, it reduces quantization error. However, this process may also introduce clipping loss, potentially harming model accuracy. To investigate this trade-off, we use AWQ as a case study to evaluate the impact of weight clipping under four bit-width settings: w4-a8, w4-a16, w8-a8 and w8-a16 on both Whisper Tiny and Whisper Base models. Table 6 reports the resulting averaged WER across all seven evaluation datasets.
| avg. WER % | w4-a8 | w4-a16 | w8-a8 | w8-a16 | ||||
| w/ clip | w/o clip | w/ clip | w/o clip | w/ clip | w/o clip | w/ clip | w/o clip | |
| Whisper Tiny | 43.67 | 21.52 | 42.53 | 17.83 | 16.76 | 14.04 | 14.80 | 12.85 |
| Whisper Base | 13.30 | 11.87 | 13.27 | 11.41 | 10.40 | 15.96 | 10.21 | 10.32 |
Our experiments reveal a clear capacity-dependent effect. For Whisper Tiny model, weight clipping consistently degrades performance across all bit-width settings. The clipping error introduced by capping weight values outweighs any reduction in quantization noise, leading to higher WER. For Whisper Base model, weight clipping begins to show measurable benefits as precision increases. At -bit weight settings, clipped models achieve modest WER reductions compared to unclipped counterparts. These findings indicate that weight clipping should be applied selectively. It can enhance quantization robustness for higher capacity models operating at higher precisions, but tend to harm smaller models, which are less resilient to additional clipping distortion.
3.2.5 Symmetric or asymmetric
Symmetric quantization with a zero offset is generally more hardware-friendly and thus preferred for deployment. However, asymmetric quantization introduces an additional grid point in the quantized space, which can improve model accuracy, especially under low bit-width settings. To quantify this effect, we compare symmetric and asymmetric weight quantization across all eight PTQ algorithms on Whisper Base and Moonshine Base models. As shown in Table 7, asymmetric quantization outperforms symmetric quantization in out of algorithms. Notably, for algorithms like SmoothQuant and AWQ, asymmetric quantization proves to be crucial to achieving acceptable WER when using -bit per-channel weight quantization. These results hightlight that, despite its slightly higher implementation complexity, asymmetric quantization often yields superior performance, making it a valuable strategy in low-bit deployment scenarios.
| Method | Whisper Base (WER %) | Moonshine Base (WER %) | ||||||
| per-group, g=64 | per-channel | per-group, g=52 | per-channel | |||||
| asym. | sym. | asym. | sym. | asym. | sym. | asym. | sym. | |
| SmoothQuant | 11.53 | 12.44 | 41.64 | 616,57 | 11.40 | 14.20 | 57.98 | 99.67 |
| AWQ | 10.95 | 11.38 | 14.12 | 50.65 | 10.39 | 10.64 | 10.75 | 11.13 |
| OmniQuant | 11.67 | 11.61 | 18.38 | 16.22 | 10.41 | 10.39 | 11.17 | 11.07 |
| GPTQ | 10.81 | 11.88 | 12.62 | 15.84 | 10.19 | 10.30 | 10.53 | 10.54 |
| RTN | 11.05 | 11.42 | 12.61 | 17.66 | 11.18 | 12.90 | 16.11 | 18.77 |
| TesseraQ | 11.14 | 11.36 | 12.41 | 17.66 | 11.16 | 12.97 | 16.09 | 18.78 |
| QUIK | 10.65 | 11.22 | 11.05 | 11.82 | 10.29 | 10.40 | 10.63 | 10.87 |
| SpQR | 10.98 | 10.93 | 11.19 | 11.14 | 10.27 | 10.15 | 10.42 | 10.51 |
3.3 Overall Comparison
Table 8 provides an overall comparisons of quantized edge-ASR models across key deployment-relevant metrics. Model size refers primarily to the total memory required to store model parameters. This has a significant impact on both latency and power consumption, particularly when weights cannot fit on-chip, resulting in frequent external memory accesses. Memory I/O captures the data movement involved in loading inputs and weights, and storing outputs for each computational operation. Since memory I/O can dominate both latency and energy usage, minimizing memory access is critical for achieving energy-efficient and low-latency inference. Bit Operations (BOPs) measures the total computational cost, taking into account both the number of MAC (multiply-accumulate) operations and data bit-width. Higher BOPs correspond to increased computing energy. Therefore, reducing bit-width not only lowers model size, but also significantly decreases the energy consumed per operation. For ASR models, the encoder is constrained by activation size, where memory I/O dominates the energy usage, while the decoder is constrained by weight size. Therefore, for low-power use cases, configurations such as w8-a8 for encoder, and w4-a16 for decoder would be considered for edge application. Based on the accuracy criteria and hardware specifications, users can choose proper models and bit-width configurations upon deployment.
| Models | # bits w/a | Encoder | Decoder | Rel. GBOPs (%) | ave. WER (%) | ||
| weight size (MB) | memory I/O (MB) | weight size (MB) | memory I/O (MB) | ||||
| Whisper tiny | 32/32 | 30.53 | 1068.35 | 118.21 | 161.55 | 100 | 12.80 |
| 4/8 | 3.82 | 263.27 | 14.78 | 25.61 | 3.13 | 14.69 | |
| 4/16 | 3.82 | 522.72 | 14.78 | 36.45 | 6.25 | 14.24 | |
| 8/8 | 7.63 | 267.09 | 29.55 | 40.39 | 6.25 | 12.94 | |
| 8/16 | 7.63 | 526.54 | 29.55 | 51.22 | 12.50 | 12.68 | |
| Whisper base | 32/32 | 79.29 | 2143.21 | 208.01 | 294.69 | 100 | 10.32 |
| 4/8 | 9.91 | 525.89 | 26.00 | 47.67 | 3.13 | 11.06 | |
| 4/16 | 9.91 | 1041.87 | 26.00 | 69.34 | 6.25 | 10.94 | |
| 8/8 | 19.82 | 535.80 | 52.00 | 73.67 | 6.25 | 10.18 | |
| 8/16 | 19.82 | 1051.78 | 52.00 | 95.34 | 12.50 | 10.17 | |
| Whisper small | 32/32 | 348.01 | 6505.61 | 614.32 | 874.36 | 100 | 8.59 |
| 4/8 | 43.50 | 1582.90 | 76.79 | 141.80 | 3.13 | 8.60 | |
| 4/16 | 43.50 | 3122.30 | 76.79 | 206.81 | 6.25 | 8.65 | |
| 8/8 | 87.00 | 1626.40 | 153.58 | 218.59 | 6.25 | 8.44 | |
| 8/16 | 87.00 | 3165.80 | 153.58 | 283.60 | 12.50 | 8.49 | |
| Moonshine tiny | 32/32 | 30.72 | 195.23 | 77.65 | 95.14 | 100 | 12.72 |
| 4/8 | 3.84 | 44.97 | 9.71 | 14.08 | 3.13 | 14.15 | |
| 4/16 | 3.84 | 86.09 | 9.71 | 18.45 | 6.25 | 13.61 | |
| 8/8 | 7.68 | 48.81 | 19.41 | 23.78 | 6.25 | 12.81 | |
| 8/16 | 7.68 | 89.93 | 19.41 | 28.16 | 12.50 | 12.73 | |
| Moonshine base | 32/32 | 80.61 | 371.40 | 165.44 | 199.13 | 100 | 9.99 |
| 4/8 | 10.08 | 82.77 | 20.68 | 29.10 | 3.13 | 10.31 | |
| 4/16 | 10.08 | 155.47 | 20.68 | 37.52 | 6.25 | 10.14 | |
| 8/8 | 20.15 | 92.85 | 41.36 | 49.78 | 6.25 | 10.04 | |
| 8/16 | 20.15 | 165.54 | 41.36 | 58.20 | 12.50 | 9.98 | |
4 Limitations
While our benchmark provides a comprehensive assessment of post-training quantization for edge-ASR models, several limitations remain. Model diversity. Our study focuses on two transformer-based ASR model families, Whisper and Moonshine, due to their compact model sizes, leading performance, and suitability for edge deployment. Future work will expand the edge-ASR models to include additional architectures (e.g., RNN-based, state space models (SSMs), Conformer variants and Canary), to broaden the applicability of our findings. Algorithm coverage. We evaluate eight state-of-the-art PTQ methods covering transformation, reconstruction and rounding-based strategies. Emerging techniques, such as rotation-based quantization, are not yet supported in our current toolkit. Incorporating these newer algorithms will enrich the benchmark and offer more effective options for aggressive model compression. Framework extensibility. Although we extend the LLM compression toolkit to support ASR models and a diverse set of quantization algorithms, many other factors remain unexplored (e.g., dynamic quantization, hardware-specific quantization strategies). Future enhancements will expand our framework to include these dimensions, enabling comprehensive support for end-to-end model optimization, conversion and deployment on a wide range of edge platforms. Addressing these limitations in future work will help guide the development of more efficient and accurate quantized ASR systems tailored for resource-constrained environments.
5 Conclusion
We present the first systematic benchmark of eight post-training quantization (PTQ) methods on two Transformer-based edge-ASR model families (Whisper and Moonshine), evaluated across seven diverse datasets. Through comprehensive analysis of critical quantization factors, including bit-width selection, ultra-low bit quantization, granularity, weight clipping, and symmetric versus asymmetric quantization, along with ablation studies on calibration data, we characterize the trade-offs between model accuracy, model size, memory footprint, I/O overhead, and computational cost in the fiexed-point domain, establishing a quantized standard for the open ASR leaderboard. Our findings show that advanced PTQ strategies enable viable quantization down to -bit weights. By building an extension to the LLM compression toolkit to support ASR architectures, we provide a workflow that enables rapid, reproducible quantization of edge-ASR models. These insights and tools serve to bridge the gap between model training and edge deployment. Future work will broaden the model zoo to include a wider range of ASR architectures, incorporate emerging quantization techniques, and integrate deployment pipelines for edge platforms, faciliating efficient and accurate quantized ASR systems on low-power, always-on edge devices.
Appendix A Architectures of edge-ASR Model Families
| Models | Whisper | Moonshine | |||
| Whisper Tiny | Whisper Base | Whisper Small | Moonshine Tiny | Moonshine Base | |
| Dimension | 384 | 512 | 768 | 288 | 416 |
| Encoder layers | 4 | 6 | 12 | 6 | 8 |
| Decoder layers | 4 | 6 | 12 | 6 | 8 |
| Attention heads | 6 | 8 | 12 | 8 | 8 |
| Encoder FFN activation | GELU | ||||
| Decoder FFN activation | GELU | GELU | GELU | SwiGLU | SwiGLU |
| Parameters in Millions | 37.8 | 72.6 | 244 | 27.1 | 61.5 |
Appendix B Quantization Mechanisms
Appendix C Dataset Description
| Dataset | Domain | Speaking Style | Test (hours) | Transcriptions |
| LibriSpeech | Audiobook | Narrated | 11 | Normalised |
| Voxpopuli | European Parliament | Oratory | 5 | Punctuated |
| TED-LIUM | TED talks | Oratory | 3 | Normalised |
| GigaSpeech | Audiobook, podcast, YouTube | Narrated, spontaneous | 40 | Punctuated |
| SPGISpeech | Financial meetings | Oratory, spontaneous | 100 | Punctuated & Cased |
| Earnings-22 | Financial meetings | Oratory, spontaneous | 5 | Punctuated & Cased |
| AMI | Meetings | Spontaneous | 9 | Punctuated & Cased |
Appendix D Model Quantization Performances
D.1 Whisper Tiny
Bits Method Whisper Tiny (WER %) AMI Earnings-22 GigaSpeech Libri clean Libri other SPGISpeech TED-Lium Voxpopuli Avg. Float Baseline 24.66 18.74 14.12 5.65 15.44 5.87 5.98 11.94 12.80 w4-a8 SmoothQuant 47.25 58.58 30.94 14.11 37.55 20.96 20.08 24.77 31.78 AWQ 38.77 37.89 20.67 10.72 20.82 12.19 12.56 18.51 21.52 OmniQuant 37.14 35.95 21.66 8.85 20.95 9.29 12.35 16.26 20.31 GPTQ 31.98 33.91 19.42 9.06 19.71 10.18 8.01 16.30 18.57 RTN 38.51 38.67 21.11 8.71 21.07 12.26 13.62 16.47 21.30 TesseraQ 30.51 25.57 17.93 8.35 20.16 9.69 9.59 14.62 17.05 QUIK 28.17 24.21 16.05 6.50 16.15 7.40 6.35 12.69 14.69 SpQR 29.63 24.93 16.71 6.70 17.99 7.71 8.56 13.79 15.75 w4-a16 SmoothQuant 37.70 42.94 22.35 8.75 20.31 12.38 12.80 17.99 21.90 AWQ 31.42 28.51 18.75 9.74 19.21 9.80 9.88 15.29 17.83 OmniQuant 33.59 26.06 19.88 7.46 17.70 8.36 8.26 13.49 16.85 GPTQ 31.65 27.94 16.92 7.07 18.26 8.20 8.00 13.60 16.46 RTN 30.84 33.28 17.59 7.69 19.78 9.79 8.70 14.54 17.78 TesseraQ 30.59 31.57 17.80 7.68 19.82 9.80 8.86 14.55 17.58 QUIK 28.21 24.56 15.95 6.35 15.98 7.37 6.36 13.05 14.73 SpQR 26.54 24.47 15.62 6.22 15.14 6.92 6.89 12.11 14.24 w8-a8 SmoothQuant 25.72 23.95 15.59 6.09 14.70 6.74 6.26 12.36 13.93 AWQ 25.37 22.04 15.03 6.54 15.74 6.71 7.75 13.13 14.04 OmniQuant 26.38 21.27 16.00 6.12 15.07 6.69 6.58 12.75 13.86 GPTQ 25.80 19.51 15.13 6.02 15.78 6.70 7.86 12.44 13.65 RTN 24.82 19.60 14.89 6.06 16.04 6.54 6.27 12.93 13.39 TesseraQ 25.47 19.95 14.33 6.25 15.03 6.26 6.21 11.75 13.16 QUIK 24.36 21.43 13.94 5.72 14.17 5.99 6.09 11.79 12.94 SpQR 26.21 22.01 14.47 5.97 15.94 6.75 7.93 13.03 14.04 w8-a16 SmoothQuant 24.36 19.35 13.97 5.69 14.64 5.87 5.99 12.18 12.76 AWQ 24.36 19.66 14.00 5.68 14.56 5.83 5.97 12.76 12.85 OmniQuant 25.99 19.17 14.52 5.68 15.41 5.90 5.93 12.14 13.09 GPTQ 24.97 20.19 14.11 5.67 14.54 5.87 5.96 12.00 12.91 RTN 24.58 18.65 13.95 5.67 15.40 5.90 5.96 12.16 12.78 TesseraQ 24.68 19.49 13.84 5.70 14.55 5.91 5.97 11.33 12.68 QUIK 23.44 19.42 14.00 5.69 15.40 5.89 5.99 11.95 12.72 SpQR 24.59 18.93 13.89 5.66 15.02 5.88 5.94 11.94 12.73
D.2 Whisper Base
Bits Method Whisper Base (WER %) AMI Earnings-22 GigaSpeech Libri. clean Libri. other SPGISpeech TED-Lium Voxpopuli Avg. Float Baseline 21.13 15.09 12.87 4.28 10.36 4.27 4.85 9.75 10.32 w4-a8 SmoothQuant 30.63 22.50 17.72 12.31 16.85 8.16 12.71 13.63 16.81 AWQ 22.77 18.95 13.35 5.16 13.13 5.01 5.35 11.23 11.87 OmniQuant 23.12 17.53 13.70 4.68 12.03 5.23 5.65 12.23 11.77 GPTQ 22.69 16.26 13.09 4.99 12.17 4.78 5.60 11.83 11.43 RTN 22.53 18.05 13.39 4.59 11.40 5.06 5.51 10.99 11.44 TesseraQ 22.03 16.79 13.46 5.20 12.35 5.06 5.46 10.89 11.40 QUIK 21.38 16.23 13.52 4.51 12.72 4.68 5.33 10.12 11.06 SpQR 22.99 15.66 13.01 4.48 11.49 4.66 5.02 11.84 11.14 w4-a16 SmoothQuant 27.13 17.82 13.76 4.68 13.30 5.41 6.54 11.27 12.49 AWQ 23.93 15.86 13.76 4.46 12.64 4.88 5.24 10.48 11.41 OmniQuant 23.06 18.00 13.89 4.64 11.99 5.05 5.39 10.71 11.59 GPTQ 23.02 16.22 13.46 4.55 11.24 4.74 5.43 10.61 11.16 RTN 23.77 16.64 13.37 4.52 11.63 4.88 5.27 10.83 11.36 TesseraQ 23.35 17.53 13.29 4.52 11.66 4.92 5.31 10.78 11.42 QUIK 21.84 16.84 13.80 4.62 12.34 4.74 5.26 10.99 11.30 SpQR 22.25 16.49 12.47 4.46 11.34 4.59 4.93 11.00 10.94 w8-a8 SmoothQuant 21.56 16.68 12.80 4.57 11.03 4.58 5.02 10.29 10.82 AWQ 21.69 15.02 12.56 4.22 10.68 4.39 4.89 9.74 10.40 OmniQuant 22.52 16.61 13.02 4.22 10.98 4.62 5.06 13.06 11.26 GPTQ 21.02 15.22 12.84 4.32 11.47 4.38 5.01 9.89 10.52 RTN 21.92 16.05 12.31 4.27 10.39 4.37 5.00 9.79 10.51 TesseraQ 20.68 16.05 12.43 4.18 11.51 4.30 4.87 10.71 10.59 QUIK 21.07 15.02 12.12 4.22 10.13 4.24 4.94 9.71 10.18 SpQR 21.35 16.37 12.50 4.25 10.66 4.44 4.93 9.84 10.54 w8-a16 SmoothQuant 21.50 15.05 12.73 4.28 10.45 4.28 4.90 10.74 10.49 AWQ 20.54 15.14 12.54 4.25 10.27 4.27 4.89 9.78 10.21 OmniQuant 21.37 15.38 12.82 4.26 10.45 4.33 4.87 9.77 10.41 GPTQ 21.58 15.04 12.74 4.29 10.63 4.26 4.89 9.79 10.40 RTN 21.15 15.05 12.59 4.25 10.46 4.28 4.89 9.71 10.30 TesseraQ 20.53 15.03 12.41 4.26 10.19 4.27 4.92 9.78 10.17 QUIK 21.13 15.08 12.69 4.28 10.16 4.23 4.91 9.78 10.28 SpQR 20.76 15.10 12.63 4.24 10.38 4.26 4.90 9.75 10.25
D.3 Whisper Small
Bits Method Whisper Small (WER %) AMI Earnings-22 GigaSpeech Libri clean Libri other SPGISpeech TED-Lium Voxpopuli Avg. Float Baseline 17.95 12.99 11.36 3.03 7.27 3.58 4.06 8.51 8.59 w4-a8 SmoothQuant 18.29 14.15 12.46 4.08 9.41 4.15 4.34 10.57 9.68 AWQ 18.51 13.14 11.10 3.18 7.92 3.82 4.38 8.52 8.82 OmniQuant 17.68 13.19 11.33 3.06 7.40 3.70 4.09 8.69 8.64 GPTQ 18.09 13.11 11.44 3.07 7.86 3.67 4.20 8.47 8.74 RTN 18.46 13.28 11.16 3.19 7.34 3.88 4.26 8.65 8.78 TesseraQ 17.96 13.40 11.00 3.10 7.78 3.84 4.15 8.61 8.73 QUIK 17.81 14.21 11.28 3.21 7.47 3.91 4.36 8.60 8.86 SpQR 18.30 12.95 11.29 3.00 7.10 3.49 4.13 8.56 8.60 w4-a16 SmoothQuant 18.03 13.52 12.11 3.33 7.81 4.17 4.36 8.64 9.00 AWQ 17.87 14.14 11.48 3.29 7.52 4.04 4.36 8.69 8.92 OmniQuant 18.59 13.85 11.54 3.21 7.28 3.73 4.20 8.98 8.92 GPTQ 18.43 13.36 11.23 3.10 7.40 3.83 4.19 8.65 8.77 RTN 17.86 13.60 11.45 3.37 7.44 4.17 4.31 8.72 8.86 TesseraQ 17.83 13.63 11.57 3.40 7.36 4.16 4.29 8.71 8.87 QUIK 17.77 14.31 11.53 3.26 7.41 3.93 4.38 8.59 8.90 SpQR 17.96 13.17 11.13 3.02 7.15 3.75 4.30 8.67 8.65 w8-a8 SmoothQuant 17.99 13.05 11.11 3.08 7.29 3.56 4.00 8.33 8.55 AWQ 17.67 13.00 11.44 3.06 7.53 3.75 4.22 8.42 8.64 OmniQuant 17.37 12.99 11.05 2.99 7.10 3.48 4.10 8.46 8.44 GPTQ 17.59 12.92 11.13 2.95 7.13 3.49 4.04 8.32 8.45 RTN 18.05 12.96 11.18 2.96 7.23 3.52 4.14 8.45 8.56 TesseraQ 17.53 12.97 11.14 2.99 7.22 3.48 3.95 8.53 8.47 QUIK 18.51 12.97 11.30 3.00 7.16 3.59 4.02 8.48 8.63 SpQR 18.02 12.96 11.14 2.97 7.43 3.54 3.98 8.30 8.54 w8-a16 SmoothQuant 17.95 13.03 11.59 3.06 7.25 3.59 4.05 8.48 8.62 AWQ 17.72 12.78 11.24 2.99 7.13 3.56 3.91 8.56 8.49 OmniQuant 17.93 13.05 11.32 3.04 7.17 3.62 4.05 8.48 8.58 GPTQ 17.93 13.00 11.44 3.08 7.24 3.63 4.05 8.51 8.61 RTN 17.95 12.96 11.36 3.06 7.55 3.64 4.10 8.48 8.64 TesseraQ 17.93 13.04 11.52 3.04 7.47 3.60 4.08 8.47 8.64 QUIK 17.91 12.98 11.41 3.06 7.46 3.63 4.05 8.49 8.62 SpQR 17.91 12.98 11.36 3.06 7.25 3.58 4.08 8.47 8.59
D.4 Moonshine Tiny
Bits Method Moonshine Tiny (WER %) AMI Earnings-22 GigaSpeech Libri clean Libri other SPGISpeech TED-Lium Voxpopuli Avg. Float Baseline 22.41 22.00 14.19 4.60 11.84 7.43 5.68 13.58 12.72 w4-a8 SmoothQuant 90.12 108.42 58.50 45.68 82.93 88.83 35.87 73.57 72.99 AWQ 26.90 25.76 16.65 5.73 14.97 9.77 7.04 15.73 15.32 OmniQuant 25.41 24.97 15.87 5.41 14.36 9.53 7.08 15.27 14.74 GPTQ 24.65 24.24 15.64 5.37 13.47 8.94 6.54 14.32 14.15 RTN 40.34 39.43 24.78 11.36 30.67 21.59 11.77 23.07 25.38 TesseraQ 30.36 31.74 19.07 6.95 17.57 12.84 8.64 16.43 17.95 QUIK 25.18 24.82 15.31 5.54 14.17 8.46 6.30 14.80 14.32 SpQR 25.07 24.42 15.64 5.28 13.70 8.69 6.47 13.98 14.15 w4-a16 SmoothQuant 88.74 107.73 53.25 41.96 77.65 80.95 34.97 73.26 69.81 AWQ 26.38 26.13 16.05 5.70 14.64 9.38 7.04 15.63 15.12 OmniQuant 24.68 24.97 15.45 5.41 13.90 9.18 7.04 15.15 14.47 GPTQ 23.79 23.27 15.03 5.09 13.01 8.22 6.37 14.12 13.61 RTN 37.72 38.64 22.13 9.37 25.39 18.12 11.88 22.55 23.23 TesseraQ 26.30 26.65 16.11 5.78 14.66 9.47 7.45 15.24 15.21 QUIK 25.03 24.64 15.19 6.00 13.59 8.35 6.21 14.88 14.24 SpQR 23.91 24.20 14.99 5.07 13.12 8.18 6.32 14.66 13.81 w8-a8 SmoothQuant 23.32 22.43 15.10 5.10 12.63 8.18 6.15 14.45 13.42 AWQ 23.25 22.34 14.71 4.82 12.32 7.94 5.78 13.78 13.12 OmniQuant 23.15 22.40 14.48 4.72 12.18 7.66 5.97 13.78 13.04 GPTQ 23.15 22.64 14.64 4.77 12.15 7.79 5.80 13.81 13.09 RTN 23.11 22.28 14.74 4.74 12.19 7.79 5.77 13.80 13.05 TesseraQ 23.93 23.55 14.91 4.82 12.79 8.05 6.04 13.93 13.50 QUIK 22.63 22.34 14.30 4.61 11.88 7.53 5.67 13.53 12.81 SpQR 23.17 22.23 14.71 4.75 12.15 7.77 5.76 13.75 13.04 w8-a16 SmoothQuant 22.47 21.99 14.19 4.60 12.00 7.46 5.68 13.81 12.78 AWQ 22.45 22.12 14.19 4.58 11.94 7.46 5.67 13.80 12.78 OmniQuant 22.34 21.96 14.21 4.61 11.90 7.44 5.63 13.76 12.73 GPTQ 22.47 22.11 14.16 4.61 11.87 7.45 5.69 13.80 12.77 RTN 22.40 22.08 14.19 4.62 11.95 7.44 5.61 13.79 12.76 TesseraQ 22.39 22.10 14.17 4.61 11.88 7.43 5.64 13.72 12.74 QUIK 22.42 22.12 14.20 4.57 11.88 7.42 5.65 13.65 12.74 SpQR 22.48 22.18 14.20 4.57 11.89 7.44 5.69 13.63 12.76
D.5 Moonshine Base
Bits Method Moonshine Base (WER %) AMI Earnings-22 GigaSpeech Libri clean Libri other SPGISpeech TED-Lium Voxpopuli Avg. Float Baseline 17.07 17.69 12.11 3.26 8.28 5.46 5.24 10.79 9.99 w4-a8 SmoothQuant 84.92 41.95 102.77 106.44 106.00 105.48 29.45 40.76 77.22 AWQ 18.47 19.72 12.84 3.77 9.46 6.05 5.83 11.74 10.98 OmniQuant 18.32 19.36 12.61 3.60 9.12 5.74 5.46 11.68 10.74 GPTQ 17.70 18.31 12.43 3.50 9.01 5.76 5.21 11.12 10.38 RTN 22.64 25.07 14.29 4.50 12.20 8.21 6.76 14.21 13.48 TesseraQ 21.41 22.05 13.49 3.95 10.67 6.95 6.04 12.04 12.08 QUIK 17.81 18.80 12.39 3.40 8.83 5.54 5.22 10.99 10.37 SpQR 17.58 17.93 12.46 3.43 9.13 5.83 5.21 10.87 10.31 w4-a16 SmoothQuant 22.43 22.94 14.22 5.55 15.77 13.78 6.33 12.57 14.20 AWQ 17.94 19.60 12.59 3.46 8.98 5.43 5.63 10.96 10.57 OmniQuant 17.88 19.16 12.50 3.51 8.87 5.66 5.48 11.25 10.54 GPTQ 17.39 18.31 12.31 3.41 8.74 5.58 5.29 11.25 10.28 RTN 21.80 24.12 13.81 4.53 11.55 7.50 6.71 13.50 12.94 TesseraQ 18.63 19.50 12.61 3.62 9.30 5.73 5.19 11.70 10.78 QUIK 17.73 18.77 12.36 3.49 8.97 5.55 5.27 11.04 10.40 SpQR 17.20 17.82 12.39 3.30 8.76 5.51 5.24 10.89 10.14 w8-a8 SmoothQuant 32.81 19.62 45.83 66.87 60.45 65.74 6.58 13.07 38.87 AWQ 17.30 17.92 12.14 3.46 8.55 5.66 5.13 11.10 10.16 OmniQuant 17.51 18.42 12.14 3.33 8.55 5.44 4.89 10.88 10.14 GPTQ 17.24 17.95 12.26 3.37 8.58 5.66 5.22 10.95 10.15 RTN 17.22 17.81 12.20 3.36 8.62 5.70 4.99 10.79 10.09 TesseraQ 17.84 18.18 12.30 3.54 8.76 5.85 5.00 10.87 10.29 QUIK 17.19 17.77 12.13 3.30 8.34 5.45 5.13 11.01 10.04 SpQR 17.26 17.82 12.24 3.36 8.52 5.73 5.14 10.88 10.12 w8-a16 SmoothQuant 16.98 17.74 12.12 3.30 8.25 5.46 5.23 10.78 9.98 AWQ 17.08 17.97 12.09 3.25 8.30 5.48 5.25 11.03 10.06 OmniQuant 17.17 17.73 12.10 3.29 8.31 5.42 5.34 11.05 10.05 GPTQ 17.10 17.71 12.10 3.27 8.29 5.47 5.17 10.83 9.99 RTN 17.11 17.62 12.10 3.29 8.27 5.50 5.28 11.07 10.03 TesseraQ 17.06 17.84 12.12 3.27 8.33 5.47 5.27 11.07 10.05 QUIK 17.12 17.69 12.10 3.28 8.29 5.47 5.31 11.05 10.04 SpQR 17.09 17.75 12.10 3.27 8.27 5.48 5.17 11.08 10.03
D.6 Runtime Cost of Quantization Algorithms
| Runtime (sec.) | SmoothQuant | AWQ | OmniQuant | GPTQ | RTN | TesseraQ | QUIK | SpQR |
| Whisper Base | 30.30 | 33409.44 | 773.13 | 85.06 | 0.78 | 331.12 | 21.79 | 69.83 |
| Moonshine Base | 32.31 | 44360.54 | 1035.37 | 99.36 | 0.76 | 978.33 | 25.28 | 90.36 |
D.7 Quantization Granularity Exeperiments
Granularity w/a Method AMI Earnings-22 GigaSpeech Libri clean Libri other SPGISpeech TED-Lium Voxpopuli Avg. Whisper Tiny w4-a16 (WER %) Float 24.66 18.74 14.12 5.65 15.44 5.87 5.98 11.94 12.80 per-channel/per-tensor AWQ 65.96 94.10 55.17 41.81 96.34 58.22 37.31 65.80 64.34 GPTQ 47.63 53.13 26.78 13.94 33.99 15.78 11.54 21.52 28.04 TesseraQ 55.04 81.12 36.94 26.08 46.40 35.19 22.32 45.54 43.58 per-group, coarse/per-tensor AWQ 41.04 46.36 23.66 8.61 22.89 13.95 10.75 18.84 23.26 GPTQ 31.52 26.25 18.23 7.74 19.19 9.04 8.25 14.48 16.84 TesseraQ 35.53 40.02 19.78 8.50 20.93 12.42 8.93 17.94 20.51 per-group, fine/per-tensor AWQ 31.42 28.51 18.75 9.74 19.21 9.80 9.88 15.29 17.83 GPTQ 31.65 27.94 16.92 7.07 18.26 8.20 8.00 13.60 16.46 TesseraQ 30.59 31.57 17.80 7.68 19.82 9.80 8.86 14.55 17.58 per-group, fine/per-token AWQ 31.51 29.60 18.61 9.74 19.26 9.79 8.77 15.26 17.82 GPTQ 33.46 28.91 17.75 7.56 16.82 8.29 8.08 12.92 16.72 TesseraQ 30.84 33.28 17.59 7.69 19.78 9.79 8.70 14.54 17.78 Whisper Base w4-a16 (WER %) Float 21.13 15.09 12.87 4.28 10.36 4.27 4.85 9.75 10.32 per-channel/per-tensor AWQ 65.54 93.77 63.01 91.68 128.17 57.10 82.33 103.53 85.64 GPTQ 31.12 26.66 16.70 5.28 15.14 6.70 7.88 13.08 15.32 TesseraQ 32.31 33.22 17.03 8.93 20.10 8.95 11.25 12.65 18.06 per-group, coarse/per-tensor AWQ 25.20 18.97 14.81 4.71 13.30 5.38 5.61 13.13 12.64 GPTQ 23.39 16.89 13.67 4.65 11.57 5.08 5.54 12.05 11.60 TesseraQ 23.28 18.09 13.32 5.11 12.45 5.26 5.54 10.91 11.75 per-group, fine/per-tensor AWQ 23.93 15.86 13.76 4.46 12.64 4.88 5.24 10.48 11.41 GPTQ 23.02 16.22 13.46 4.55 11.24 4.74 5.43 10.61 11.16 TesseraQ 23.35 17.53 13.29 4.52 11.66 4.92 5.31 10.78 11.42 per-group, fine/per-token AWQ 23.68 15.87 14.01 4.48 12.24 4.94 5.25 10.56 11.38 GPTQ 25.49 16.68 13.56 4.64 12.03 4.97 5.44 12.25 11.88 TesseraQ 23.77 16.64 13.37 4.52 11.63 4.88 5.27 10.83 11.36 Whisper Small w4-a16 (WER %) Float 17.95 12.99 11.36 3.03 7.27 3.58 4.06 8.51 8.59 per-channel/per-tensor AWQ 20.59 14.46 12.23 4.01 9.52 4.70 4.45 9.59 9.94 GPTQ 18.32 14.61 12.14 3.36 8.24 4.24 4.23 9.01 9.27 TesseraQ 19.18 13.59 11.77 3.67 9.24 4.36 4.47 9.13 9.42 per-group, coarse/per-tensor AWQ 18.05 13.39 11.38 3.16 7.65 4.21 4.40 8.68 8.86 GPTQ 17.92 13.36 11.68 3.17 7.54 3.80 4.21 8.78 8.81 TesseraQ 17.88 13.87 11.56 3.49 7.53 4.23 4.40 8.74 8.96 per-group, fine/per-tensor AWQ 17.87 14.14 11.48 3.29 7.52 4.04 4.36 8.69 8.92 GPTQ 18.43 13.36 11.23 3.10 7.40 3.83 4.19 8.65 8.77 TesseraQ 17.83 13.63 11.57 3.40 7.36 4.16 4.29 8.71 8.87 per-group, fine/per-token AWQ 17.79 14.16 11.47 3.35 7.62 4.22 4.37 8.67 8.96 GPTQ 18.99 13.49 11.51 3.56 7.36 3.80 4.09 8.79 8.95 TesseraQ 17.86 13.60 11.45 3.37 7.44 4.17 4.31 8.72 8.86 Moonshine Tiny w4-a16 (WER %) Float 22.41 22.00 14.19 4.60 11.84 7.43 5.68 13.58 12.72 per-channel/per-tensor AWQ 27.25 27.08 16.53 6.02 15.37 9.99 7.60 15.51 15.67 GPTQ 25.70 25.22 15.63 5.95 14.16 8.60 6.75 14.80 14.60 TesseraQ 27.79 27.94 17.10 6.07 17.04 9.96 7.95 16.27 16.26 per-group, coarse/per-tensor AWQ 25.95 26.11 16.14 5.86 15.12 9.37 7.22 15.62 15.17 GPTQ 24.39 23.99 15.31 5.14 13.41 8.27 6.53 14.24 13.91 TesseraQ 26.98 27.05 16.68 5.88 15.34 9.96 7.41 15.92 15.65 per-group, fine/per-tensor AWQ 26.38 26.13 16.05 5.70 14.64 9.38 7.04 15.63 15.12 GPTQ 23.79 23.27 15.03 5.09 13.01 8.22 6.37 14.12 13.61 TesseraQ 26.30 26.65 16.11 5.78 14.66 9.47 7.45 15.24 15.21 per-group, fine/per-token AWQ 26.10 25.85 15.84 5.48 14.51 9.22 7.00 15.57 14.95 GPTQ 23.86 23.35 14.94 5.07 13.00 8.36 6.35 14.42 13.67 TesseraQ 26.42 26.71 16.17 5.76 14.68 9.43 7.49 15.26 15.24 Moonshine Base w4-a16 (WER %) Float 17.07 17.69 12.11 3.26 8.28 5.46 5.24 10.79 9.99 per-channel/per-tensor AWQ 19.03 20.65 12.97 3.69 9.56 5.93 5.40 12.06 11.16 GPTQ 17.93 19.55 12.61 3.52 9.02 6.05 5.79 11.44 10.74 TesseraQ 19.65 20.59 13.11 3.86 9.90 6.27 5.66 12.41 11.43 per-group, coarse/per-tensor AWQ 18.18 20.16 12.60 3.59 9.19 5.86 5.21 11.53 10.79 GPTQ 17.93 18.78 12.60 3.43 8.89 5.70 5.31 11.38 10.50 TesseraQ 18.83 20.77 12.84 3.58 9.75 6.22 5.35 12.25 11.20 per-group, fine/per-tensor AWQ 17.94 19.60 12.59 3.46 8.98 5.43 5.63 10.96 10.57 GPTQ 17.39 18.31 12.31 3.41 8.74 5.58 5.29 11.25 10.28 TesseraQ 18.63 19.50 12.61 3.62 9.30 5.73 5.19 11.70 10.78 per-group, fine/per-token AWQ 18.07 19.39 12.62 3.46 8.89 5.54 5.62 11.57 10.64 GPTQ 17.60 18.42 12.19 3.45 8.67 5.47 5.45 11.14 10.30 TesseraQ 18.54 19.34 12.63 3.65 9.36 5.73 5.18 11.69 10.77
D.8 Weight Clipping Exeperiments
Bits Method AMI Earnings-22 GigaSpeech Libri clean Libri other SPGISpeech TED-Lium Voxpopuli Avg. Whisper Tiny (WER %) Float 24.66 18.74 14.12 5.65 15.44 5.87 5.98 11.94 12.80 w4-a8 w/ clip 107.70 83.45 48.37 11.59 44.29 19.48 10.89 23.61 43.67 w/o clip 38.77 37.89 20.67 10.72 20.82 12.19 12.56 18.51 21.52 w4-a16 w/ clip 115.78 76.79 42.04 10.08 42.67 14.63 13.97 24.29 42.53 w/o clip 31.42 28.51 18.75 9.74 19.21 9.80 9.88 15.29 17.83 w8-a8 w/ clip 32.74 31.41 17.70 6.73 18.02 7.22 8.23 12.01 16.76 w/o clip 25.37 22.04 15.03 6.54 15.74 6.71 7.75 13.13 14.04 w8-a16 w/ clip 36.31 22.06 15.30 6.18 14.42 6.13 6.16 11.83 14.80 w/o clip 24.36 19.66 14.00 5.68 14.56 5.83 5.97 12.76 12.85 Whisper Base (WER %) Float 21.13 15.09 12.87 4.28 10.36 4.27 4.85 9.75 10.32 w4-a8 w/ clip 26.86 21.73 16.08 5.26 13.28 5.69 5.71 11.82 13.30 w/o clip 22.77 18.95 13.35 5.16 13.13 5.01 5.35 11.23 11.87 w4-a16 w/ clip 26.77 20.62 15.95 5.67 14.14 5.53 6.11 11.35 13.27 w/o clip 23.93 15.86 13.76 4.46 12.64 4.88 5.24 10.48 11.41 w8-a8 w/ clip 21.69 15.02 12.56 4.22 10.68 4.39 4.89 9.74 10.40 w/o clip 28.50 19.35 17.48 9.90 19.37 7.42 9.76 15.91 15.96 w8-a16 w/ clip 20.54 15.14 12.54 4.25 10.27 4.27 4.89 9.78 10.21 w/o clip 21.14 15.08 12.74 4.28 10.42 4.23 4.90 9.75 10.32
D.9 Symmetric or Asymmetric Quantization Exeperiments
Granularity Method Config. AMI Earnings-22 GigaSpeech Libri clean Libri other SPGISpeech TED-Lium Voxpopuli Avg. Whisper Base (WER %) Float 21.13 15.09 12.87 4.28 10.36 4.27 4.85 9.75 10.32 per-channel SmoothQuant sym. 560.13 548.78 554.81 636.43 702.38 702.82 614.35 612.87 616.57 asym. 34.14 55.86 35.83 46.82 47.00 27.18 47.05 39.26 41.64 AWQ sym. 55.48 65.68 38.30 42.84 75.90 28.94 37.34 60.71 50.65 asym. 25.44 22.36 14.87 5.25 13.74 6.76 7.01 17.52 14.12 OmniQuant sym. 31.65 25.61 19.18 6.46 15.53 7.78 6.43 17.13 16.22 asym. 36.16 32.18 20.57 5.87 18.09 8.82 8.71 16.62 18.38 GPTQ sym. 27.38 25.61 16.53 7.59 17.56 7.42 5.87 18.80 15.84 asym. 24.18 21.68 14.11 4.74 13.17 5.25 5.64 12.20 12.62 RTN sym. 29.85 32.33 17.24 9.20 19.78 8.91 11.16 12.81 17.66 asym. 24.33 19.64 13.97 5.40 13.47 5.78 5.69 12.63 12.61 TesseraQ sym. 29.85 32.33 17.24 9.20 19.78 8.91 11.16 12.81 17.66 asym. 23.84 18.34 14.09 5.29 13.73 5.65 5.76 12.59 12.41 QUIK sym. 22.43 16.99 13.62 5.15 13.24 5.53 5.97 11.63 11.82 asym. 21.46 16.84 13.26 4.58 11.18 4.75 5.28 11.02 11.05 SpQR sym. 22.23 16.79 12.94 4.58 12.17 4.81 5.35 10.24 11.14 asym. 22.61 17.14 12.94 4.53 11.71 4.90 5.45 10.19 11.19 per-group SmoothQuant sym. 27.35 17.85 14.00 5.04 12.52 5.34 6.54 10.92 12.44 asym. 23.63 18.46 13.56 4.67 11.37 4.95 5.36 10.24 11.53 AWQ sym. 23.68 15.87 14.01 4.48 12.24 4.94 5.25 10.56 11.38 asym. 22.28 16.79 12.76 4.94 10.98 4.60 5.12 10.15 10.95 OmniQuant sym. 23.11 18.65 13.79 4.65 11.53 4.92 5.46 10.73 11.61 asym. 23.35 16.45 14.37 4.55 12.82 4.97 5.42 11.41 11.67 GPTQ sym. 25.49 16.68 13.56 4.64 12.03 4.97 5.44 12.25 11.88 asym. 22.34 15.72 13.34 4.43 10.87 4.49 5.12 10.15 10.81 RTN sym. 23.35 17.53 13.29 4.52 11.66 4.92 5.31 10.78 11.42 asym. 21.70 16.35 13.24 4.99 12.15 4.69 5.10 10.16 11.05 TesseraQ sym. 23.77 16.64 13.37 4.52 11.63 4.88 5.27 10.83 11.36 asym. 21.90 16.96 13.25 5.01 12.20 4.58 5.15 10.06 11.14 QUIK sym. 21.80 17.28 13.76 4.61 12.08 4.76 5.26 10.22 11.22 asym. 20.40 16.03 12.67 4.46 11.36 4.44 4.98 10.86 10.65 SpQR sym. 21.62 16.39 12.88 4.42 11.32 4.57 5.24 10.97 10.93 asym. 22.40 15.54 13.21 4.71 11.08 4.49 5.04 11.35 10.98 Moonshine Base (WER %) Float 17.07 17.69 12.11 3.26 8.28 5.46 5.24 10.79 9.99 per-channel SmoothQuant sym. 78.60 106.55 94.42 106.85 102.91 95.70 98.95 113.42 99.67 asym. 42.40 59.07 52.15 72.20 70.41 48.03 54.19 65.36 57.98 AWQ sym. 19.06 20.38 12.96 3.60 9.78 6.08 5.30 11.87 11.13 asym. 18.77 19.28 12.68 3.51 9.29 5.64 5.38 11.49 10.75 OmniQuant sym. 19.41 20.27 12.73 3.78 9.53 5.81 5.38 11.62 11.07 asym. 19.60 20.62 12.94 3.75 9.53 5.82 5.25 11.83 11.17 GPTQ sym. 17.83 18.82 12.59 3.48 8.89 5.92 5.53 11.26 10.54 asym. 17.84 18.77 12.58 3.52 8.99 5.78 5.52 11.24 10.53 RTN sym. 35.65 32.53 17.50 6.23 19.64 11.85 8.61 18.15 18.77 asym. 29.19 29.84 16.07 4.78 14.61 10.15 6.83 17.37 16.11 TesseraQ sym. 35.64 32.65 17.52 6.25 19.56 11.88 8.53 18.18 18.78 asym. 29.02 29.72 16.04 4.75 14.82 9.85 6.92 17.63 16.09 QUIK sym. 19.06 19.62 12.63 3.59 9.52 5.79 5.30 11.43 10.87 asym. 18.58 18.59 12.53 3.43 9.15 6.07 5.25 11.46 10.63 SpQR sym. 18.11 18.50 12.59 3.39 8.97 5.76 5.23 11.51 10.51 asym. 17.87 18.28 12.46 3.47 8.76 6.08 5.30 11.15 10.42 per-group SmoothQuant sym. 22.44 22.92 14.20 5.51 15.88 13.76 6.34 12.58 14.20 asym. 20.41 19.18 12.85 4.15 10.51 7.06 5.03 11.97 11.40 AWQ sym. 18.07 19.39 12.62 3.46 8.89 5.54 5.62 11.57 10.64 asym. 17.94 18.80 12.47 3.37 8.65 5.50 5.50 10.89 10.39 OmniQuant sym. 17.81 18.41 12.44 3.45 9.04 5.64 5.21 11.11 10.39 asym. 17.88 18.41 12.41 3.49 8.81 5.54 5.60 11.16 10.41 GPTQ sym. 17.60 18.42 12.19 3.45 8.67 5.47 5.45 11.14 10.30 asym. 17.28 18.20 12.24 3.36 8.52 5.59 5.17 11.15 10.19 RTN sym. 21.78 23.95 13.81 4.52 11.50 7.50 6.60 13.52 12.90 asym. 19.05 20.33 12.71 3.66 9.58 6.07 5.33 12.71 11.18 TesseraQ sym. 21.88 24.17 13.84 4.58 11.55 7.52 6.69 13.54 12.97 asym. 18.89 20.38 12.66 3.52 9.58 6.17 5.18 12.93 11.16 QUIK sym. 17.74 18.80 12.37 3.48 8.96 5.54 5.25 11.05 10.40 asym. 17.88 18.47 12.32 3.45 8.46 5.35 5.26 11.12 10.29 SpQR sym. 17.31 17.80 12.37 3.35 8.71 5.47 5.28 10.91 10.15 asym. 17.56 18.28 12.24 3.35 8.55 5.62 5.45 11.07 10.27
Appendix E Ablation Study
To evaluate how calibration data affects quantization performance, we vary both the source and size of the calibration set. Unless otherwise specified, experiments use -bit per-group weight and -bit per-tensor activation quantization.
Calibration data sources. We compare two calibration sources: A held-out subset of English utterances from Mozilla Common Voice (our default), and a subset of the same size sampled from LibriSpeech Dev. set. Across all eight PTQ algorithms and both models (Whisper Base and Moonshine Base), switching between these calibration sources results in WER variations of less than , suggesting that most PTQ methods are insensitive to the specific calibration speech corpus, under a good number of sample sizes.
| avg. WER % | Whisper base | Moonshine base | ||
| Calib. Data | Com. Voice | Libri. Dev | Com. Voice | Libri. Dev |
| SmoothQuant | 12.49 | 12.40 (-0.09) | 14.20 | 12.38 (-1.82) |
| AWQ | 11.41 | 11.80 (+0.39) | 10.57 | 10.67 (+0.10) |
| OmniQuant | 11.59 | 11.72 (+0.13) | 10.54 | 10.53 (-0.01) |
| GPTQ | 11.16 | 11.62 (+0.46) | 10.28 | 10.40 (+0.12) |
| RTN | – | – | – | – |
| TesseraQ | 11.42 | 11.37 (-0.05) | 12.98 | 12.94 (-0.04) |
| QUIK | 11.30 | 11.10 (-0.20) | 10.40 | 10.33 (-0.07) |
| SpQR | 10.94 | 10.85 (-0.09) | 10.14 | 10.17 (+0.03) |
| avg. WER % | Whisper base | Moonshine base | ||||
| Calib. Samples | 128 | 256 | 512 | 128 | 256 | 512 |
| SmoothQuant | 12.25 | 12.49 | 12.78 | 14.61 | 14.20 | 13.57 |
| AWQ | 11.94 | 11.41 | 97.04 | 10.60 | 10.57 | 18.23 |
| OmniQuant | 11.86 | 11.59 | 12.21 | 10.47 | 10.54 | 10.53 |
| GPTQ | 11.23 | 11.16 | 11.27 | 10.36 | 10.28 | 10.33 |
| RTN | – | – | – | – | – | – |
| TesseraQ | 11.37 | 11.42 | 11.37 | 12.94 | 12.98 | 12.94 |
| QUIK | 10.92 | 11.30 | 10.78 | 10.38 | 10.40 | 10.36 |
| SpQR | 10.84 | 10.94 | 10.81 | 10.25 | 10.14 | 10.29 |
Calibration data sizes. Using Mozilla Common Voice, we further vary the calibration set size across , and samples. Increasing the set from to samples causes minimal changes in WER across all algorthms. However, when increased to samples, AWQ exhibits a sharp performance drop. From the block-wise activation distribution shown in Figure 4, a possible reason is that extreme activation outliers dominating the scale computation during -bit weight quantization, which destroys quantization representation for weights. In contrast, other algorithms including reconstruction-based and rounding optimizations, remain stable at samples. These results suggest that, while most PTQ techniques are generally robust to calibration speech corpus, algorithms sensitive to activation outliers (e.g., AWQ) may require careful tuning of calibration set size to avoid performance degradation.